1Is the sequence geometric? If so, identify the common ratio.
1/4, 3/16, 9/64, 27/256, 81/1024.
a) yes: 1/3
b) yes: 3/4
c) not geometric
d) yes: 2/9

Answer by fcabanski(1391)

Since this is multiple choice, check each answer from term to term. A term times the common ratio equals the next term.


A: 1/3: 1/4 * 1/3 = 1/12. That's not the second term. Eliminate answer A.


B: 3/4: 1/3 * 3/4 = 3/16. It works as a ratio between the first and second term.


3/15 * 3/4 = 9/64. Good for third term.


9/64 * 3/4 = 27/256. Good for fourth term.


27/256 * 3/4 = 81/1024. This ratio works for all terms, so the answer is b.


You can also divide each term by the previous term to find the common ratio. If the common ratio between terms is the same, then it's a geometric sequence, and you have the common ratio.


81/1024 divided by 27/256 = 81/1024 * 256/24 = 3/4. Do the same for each pair terms to find that the common ratio between all terms is 3/4.